OpenStudy (anonymous):

How do you put a hyperbolic arctan into a TI 83 calculator?

zepdrix (zepdrix):

Umm so \(\tanh x\) can be written with exponentials. So I think \(arctanh x\) can be written with logs. I think this is what you get,\[\large arctanh x=\frac{1}{2}\frac{\ln(1+x)}{\ln(1-x)}\]

zepdrix (zepdrix):

You might wanna google it just to make sure, I'm a little rusty on identities D: Putting that into your calculator should get you what you need c:

OpenStudy (anonymous):

ahhhh i finally found it online thru the manuel... took awhile... but just in case you wondered, there is a tanh(-1), button under the catalog selection of the calculator... thanks for looking.

OpenStudy (anonymous):

i also found that formula you just mentioned, which worked too... but it was .5 ln (1+r/1-r)

zepdrix (zepdrix):

TI's have the hyperbolic functions? nice!